Debbie traveled by boat 5 miles upstream to
fish in her favorite spot. Because of the 4-mph current, it
took her 20 minutes longer to get there than to return. How
fast will her boat go in still water?

Answer by mananth(16946)

Debbie traveled by boat 5 miles upstream to
fish in her favorite spot. Because of the 4-mph current, it
took her 20 minutes longer to get there than to return. How
fast will her boat go in still water?
...
let still water speed = x mph
downstream speed = x+4 mph
upstream speed = x-4 mph
time upstream = 5/(x-4)
time downstream = x/(x+4)
time upstream - time donstream = 1/3 hours
5/(x-4)-5/(x+4)=1/3
5(x+4)-5(x-4)/(x+4)(x-4)=1/3
5x+20-5x+20=(x+4)(x-4)/3
40=(x^2-16)/3
120 = x^2-16
x^2=136
x = 11.66 mph
...
m.ananth@hotmail.ca